OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4). −11 21 2.25 11

OpenStudy (anonymous):

How do you think you should proceed?

jimthompson5910 (jim_thompson5910):

g(x) = −4x + 5 g(-4) = −4(-4) + 5 g(-4) = 16 + 5 g(-4) = 21

OpenStudy (anonymous):

thank you for showing the steps :)